OpenStudy (anonymous):

There are 6 pounds of potatoes. Tammy needs eight-ninths of a pound of potatoes to make one serving of a dish. How many servings can she make? four-twenty-sevenths serving nine forty-eighths serving five and one-third servings six and three-fourths servings

OpenStudy (mathstudent55):

This is a division problem. You need to divide 6 by 8/9. Do you know how to divide fractions?

OpenStudy (anonymous):

@zaibali.qasmi im so happy ur her

OpenStudy (anonymous):

see..you just have to divide them... total amount=6pounds required for one=8/9 to find how many he can make with 6 pounds...just divide \[\frac{ 6 }{ \frac{ 8 }{ 9 } }=6\times \frac{ 9 }{ 8 }=\frac{ 54 }{ 8 }\] nnow divide it and get result..

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

mind giving me the answer im in a hurry

OpenStudy (anonymous):

which thing not getting..?

OpenStudy (anonymous):

answer plz

OpenStudy (anonymous):

six and three forth...see how.. \[6 \times \frac{ 3 }{ 4 }\] now multiply and divide by 2 it will become.. \[=6 \times \frac{ 6 }{ 8 }=\frac{ 54 }{ 8 }\] as i explain in above coment

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

read my last coment starting words...

OpenStudy (anonymous):

answer is option D...

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

you are warmly welcomme..

OpenStudy (anonymous):

hehe